this conical glass holds 150ml. its top radius is 4 cm find its depth
Answer by rothauserc(4718)

convert 150ml to cubic cm
:
1 ml is 1 cubic cm, therefore 150 ml = 150 cubic cm
:
150 = 1/3 × pi × r^2 × h
:
Note conical means cone shaped
:
150 = (1/3) * (22/7) * 4^2 * d, where d is depth
:
150 = 352d / 21
:
352d = 150 * 21
:
d = 8.94886 approx 8.9
:
**************************************
depth of glass is approximately 8.9 cm
**************************************
: